OpenStudy (anonymous):

Lost in step 3: 2. Apply the formula of a right circular cylinder (V = r2h) to find the volume of the object. 3. Rewrite the formula using the variable x for the radius. Substitute the value of the volume found in step 2 for V and express the height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the radius, then the expression for the height will be (x + 4). Diameter: 3in Height: 4.5in Volume: ≈31.8in

OpenStudy (kinggeorge):

I think what it's saying is to let height be the radius (x) plus some value. Since your height is 4.5, and your diameter is 3, the radius is 1.5, so height = x+3. Also, V~31.8, so you would rewrite your equation as V=x2(x+3)

OpenStudy (anonymous):

ooops thinking SA

OpenStudy (anonymous):

v=pi r^2 h

OpenStudy (anonymous):

did you forget pi ?

OpenStudy (anonymous):

Yes, pi should be there. :)

OpenStudy (kinggeorge):

I'm a little confused by your original formula though. I though that \[V=\pi r^2 h\]Like mathg8 said.

OpenStudy (anonymous):

Sorry, copy and pasted. Didn't realize that the pi symbol didn't transfer.

OpenStudy (kinggeorge):

So then, using what I said earlier, \[V=\pi x^2 (x+3)\]

OpenStudy (anonymous):

Now I: Simplify the equation and write it in standard form. If the equation contains decimals, multiply each term by a constant that will make all coefficients integers. Example: 0 = x3 + 2.75x2 – 6.25 If this equation is multiplied by 4, the decimals will become integers. 4(0 = x3 + 2.75x2 – 6.25) 0 = 4x3 + 11x2 – 25

OpenStudy (kinggeorge):

Since \(V \sim 31.8\) you have \(31.8=\pi x^3+\pi 3x^2\). Here, \(31.8\) is the only decimal, so multiply by \(5\) to get rid of the decimal.

OpenStudy (anonymous):

159=5pix^2+5pi3x^2 ?

OpenStudy (kinggeorge):

So you get \(0 \sim 5\pi x^3 + 15\pi x^2 -159\). Now, the only problem I'm seeing, is that \(\pi\) isn't an integer, but there's no way to get rid of it. However, if you kept the original value of \(V\) in terms of \(\pi\), we could get rid of pi.

OpenStudy (kinggeorge):

So, assuming that, we can use the original diameter and height to find that \[V=10.125 \pi = \pi x^3 +3 \pi x^2\]This is much better since we can now divide by \(\pi\) and multiply by \(8\) since \(0.125 = {1/8}\). Finally, that means that \[0=8x^3+24x^2-81\]

OpenStudy (anonymous):

Thank you! I'm taking this class online, and have e-mailed my teacher asking for help multiple times. They don't seem as though they are willing to help, as they have never responded. Lol. :(

OpenStudy (kinggeorge):

You're very welcome.

OpenStudy (kinggeorge):

Which class is this?

OpenStudy (anonymous):

Algebra II

OpenStudy (kinggeorge):

I've always found algebra interesting.

OpenStudy (anonymous):

Math is not a skill of mine, so I would have to disagree! Thank you again. :)

OpenStudy (kinggeorge):

You're welcome. Feel free to ask any more questions.